the radius of the wire in a coaxial cable is 0.65mm and the inner radius of the co-axial conducting cylinder is 1.45mm.asssuming that there is vacuum between the wire and the cylinder,calculate the capacitance of a 1.5m length of the cable.

Grade:12th pass

1 Answers

Eshan
5 years ago
Capacitance of a cylindrical capacitor is given as

C=\dfrac{2\pi k\epsilon_0L}{ln(\dfrac{r_2}{r_1})}
=\dfrac{2\pi\epsilon_0(1.5)}{ln(\dfrac{1.45}{0.65})}=1.04\times 10^{-10}F